Dublin University Rifle Club

Version Date : April 24, 2002 Filename : Status :

Dublin University Rifle Club Secretary’s Report

Mark Dennehy Chief Range Officer

April 24, 2002

1 The last four years

I’m leaving the committee this year after three and a half years serving on it, mainly because I believe noone should stay on a committee too long. Looking back on the last few years, I’m quite pleased with what the committees from the last few years have accomplished. We’ve taken the club from the 1995-6 slump where we had about thirty people actually shooting to the current high of 436 members, nearly a hundred active, the ladders full of cards and the colours team about to stomp all over UCD. Our financial situation isn’t rich, but it’s a great deal better than when we found it. The standard of shooting in the club has risen enormously, mainly thanks to Steven and John, and our good name is now well established.

I mention these achievements for the incoming committee – when we started, we set ourselves the goals of bringing the membership back up, increasing our attendance at competitions, fixing the financial problems the club had and getting our members more active in the club again. We had been in the club during the pre-95 period and so had a good idea of what we were aiming at. The incoming commitee should pick goals for themselves to accomplish and let these decide what they do - it worked wonders for the current committee, and it’s really the best way to tackle the job of running the club.

2 New Air Rifle Classification Scheme

NTSA Air Rifle coordinator Geoff Cooney proposed a new classification system for Irish shooters this year. Discussions took place between him, us and UCD. The new system would see the introduction of sub-classes to the current classes and a general all-round renaming policy. To promote the sense of advancement through

3 Archival of Club Documents

the ranks, it was thought that each of the four current classes would be split into three sub-classes. You could move from sub-class to sub-class in internal club events, but only from class to class in external events. We would use our ladder to decide which sub-class a shooter was in. The idea did mean a bit more work for stats people, and there was the spectre of needing forty-odd prizes (3 places × 3 sub-classes × 4 classes plus prizes for Junior and Ladies classes). In practise, this wouldn’t happen except in really rich clubs , with prizes begin given out for the four classes only. The sceme also involved expanding the Junior category to includ age brackets (u14, u16, u18) with different courses of fire for each (20 shots for u14, 30 for u16 and 40 for u18), to prevent spinal problems in the very young juniors, a serious problem as I, or Ray, or Steven, or any other old person can testify to!

The scores for the classes were set as follows:

Adult Male

Novice Intermediate Senior Master
Grade 3 Grade 2 Grade 1 Grade 3 Grade 2 Grade 1 Grade 3 Grade 2 Grade 1 Grade 3 Grade 2 Grade 1
60*5  60*5.5  60*6  60*6.5  60*7  60*7.5  60*8  60*8.5  60*9  60*9.5 60*9.6 60*9.8
300  330  360  390  420  450  480  510  540  570  576  588

Adult Female

Novice Intermediate Senior Master
Grade 3 Grade 2 Grade 1 Grade 3 Grade 2 Grade 1 Grade 3 Grade 2 Grade 1 Grade 3 Grade 2 Grade 1
40*5  40*5.5  40*6  40*6.5  40*7  40*7.5  40*8  40*8.5  40*9  40*9.5 40*9.6 40*9.8
200  220  240  260  280  300  320  360  380  384  392

Note that the ladies now shoot a 40-shot match, according to international standards. This has not been the case in Ireland up to now, as there were never enough women shooters to justify it. However, the change should be made as our international performances show a deleterious effect on Irish shooters’ results because they have to adapt to a shorter match.

The system was not used this year - it is due to be phased in over the summer and to be in full swing by next October. The four Air Rifle competitions this summer are designed to break-in and debug the system, where it will be run in parallel to the old 4-class system.

3 Archival of Club Documents

The “Markive” of old club documents on the range was put together at the end of last summer. I found a lot of interesting stuff in these documents - the original minutes book for the commitee, and the old Match Secretary records and so forth. However, the records are physically deteriorating due to the environment. The old library has agreed to archive the material for us – the next secretary should arrange to transfer the material to them, as they’re eager to have it and we have been assured that our records will remain confidential.

4 Freedom of Information Act

This year saw the FOI act brought into force in college. So far we have had no requests under it, but our records are now better kept, so we could if necessary respond to one.

5 NTSA AGM

The NTSA AGM took place on April 6 this year. This is important to us as the NTSA makes decisions as to the sport in the country as a whole - like the new Air classification system for example. They also organise

6 Alumni Lists

the National Championships, National Squad, training and coaching weekends and decide who can represent Ireland abroad. This year we have myself on committee again as PRO. Complaints were put forward by DURC via John Keeney this year regarding the smallbore classification system which has 50m shooters being classified with 25yd scores, and with the NTSA Calendar where our competitions were shoved around to make room for the national squad training weekends. We’ve also just issued a complaint as the NTSA scheduled a shoot on the same day as our Intervarsity match, which is prejudicial towards our shooters. Nothing on this has yet been resolved however.

6 Alumni Lists

The Alumni office has suggested that we submit details of all DURC shooters/ROs/Committee members who leave college each year to facilitate reunions and so forth. Having seen first-hand how hard it is to find exshooters once they leave college (though we’re currently seeing a reemergence of some of our more prominent ex-shooters like Conor Tighe and Simon Kennedy), I believe it’s an excellent idea. There are some practical questions related to figuring out who is leaving, whether they want to stay in touch, and general paperwork work, but on the whole, this is worth the effort.

7 Government Grant woes

We tried applying for a government Department of Sport grant this year for capital equipment. Sadly, we were turned down on the basis that we were the Department of Education’s responsibility - of course they say we’re the Dept.of Sport’s baliwick...

The next committee should lobby to get into the scheme for next year. 76 million euro was given out this year - we should be trying to get some of that as if we had, we could have bought two rifles this year instead of one!

8 Club Sponsorship

Efforts to get sponsorship for the club this year proved to be a mixed bag. Government grants were ruled out as I outlined above, but private and corporate sponsorship was also sought. Eircom turned us down, as did Bank of Ireland, AIB, Eircell, Esat, and Meteor. We also appealed to target shooting equipment companies, asking for sponsorship of events instead of the club, and here we had more luck, with Hamerelli giving us a box of caps for the DURC 10m Open and Shilen giving us a voucher for a barrel for the 50m Open. Simon Kennedy also donated privatly to the prize fund for the 10m Open, and now has a shield in his name.

Next year’s committee has to start looking for sponsorship much earlier - as in right after exam time. Eircom and the others decide at the start of the year who they’ll sponsor - we need to hit them with a list of our achievements in the last year, articles in the press about us, our long safety record and so forth, and we need to be quite proactive about it - until you talk to them, and not just write, you can’t write them off.

9 Management Problems

This is our biggest problem in DURC. We’re a volunteer organisation. We’re also the biggest club in college, catering to 436 members and doing so with minimal resources in a very demanding sport. This means that the committee are being swamped in day-to-day running of the club and have little time left for projects either major or minor. Most of us have academic responsibilies, exams, courses to teach, PhD’s to write or jobs to hold down as well as train; and that’s before we do our DURC committee work. There’s a long history in DURC of committee members failing courses and leaving college or repeating due to their workload in committee demanding too much of them. A new scheme is needed.

9 Management Problems

Basically, the problem is that we currently manage the club on a “line” management philosophy - ie. day-to-day running of a continous process. We have no facility for individual projects, such as introducing 3P Air Rifle, or construction of things for the range. In the past, non-committee people like Keith have done all the DIY. Well appreciated, and the club thanks Keith for this, but the club needs a mechanism to be able to do things like this themselves. We need, in short, a pool of people who are not on committee to do specific projects - time-limited, single-task jobs, like “build the target holders for 3P air rifle”. I urge the incoming committee to develop such a pool and to codify it and write down somewhere how they want it to work.

Otherwise, the incoming committee will get as stressed with workload as this year’s outgoing committee have gotten.